# Deep Spectral Prior

### v1

@@ Abstract
We introduce Deep Spectral Prior (DSP), a new formulation of Deep Image Prior (DIP) that redefines image reconstruction as a frequency-domain alignment problem. Unlike traditional DIP, which relies on pixel-wise loss and early stopping to mitigate overfitting, DSP directly matches Fourier coefficients between the network output and observed measurements. This shift introduces an explicit inductive bias towards spectral coherence, aligning with the known frequency structure of images and the spectral bias of convolutional neural networks. We provide a rigorous theoretical framework demonstrating that DSP acts as an implicit spectral regulariser, suppressing high-frequency noise by design and eliminating the need for early stopping. Our analysis spans four core dimensions establishing smooth convergence dynamics, local stability, and favourable bias–variance tradeoffs. We further show that DSP naturally projects reconstructions onto a frequency-consistent manifold, enhancing interpretability and robustness. These theoretical guarantees are supported by empirical results across denoising, inpainting, and super-resolution tasks, where DSP consistently outperforms classical DIP and other unsupervised baselines.
@@ 1 Introduction
Inverse problems lie at the heart of image restoration and reconstruction, encompassing tasks such as denoising, inpainting, super-resolution, and deblurring. Traditionally approached through handcrafted priors and variational optimisation, these problems have seen a paradigm shift with the advent of deep learning, with single-image techniques effectively leverage data essakine2025where ; cheng2024singleshot ; cheng2024you . Among the most intriguing developments in this space is the Deep Image Prior (DIP) ulyanov2018deep , which showed that the structure of an untrained convolutional neural network (CNN) alone can act as a powerful prior. By optimising a randomly initialised network to match a single degraded observation, DIP enables high-quality reconstructions—without requiring any training data. This surprising result has made DIP a widely adopted framework across diverse domains, including hyperspectral rasti2021undip , microscopy vu2021deep ; zhou2020diffraction , compressed sensing van2018compressed , and medical imaging gong2018pet ; yoo2021time ; barbano2022educated ; liang2025analysis ; liang2025sequential .
@@
Despite its empirical success, DIP suffers from two key limitations that hinder its practical effectiveness. First, DIP is notoriously prone to overfitting. Due to the lack of explicit regularisation, the network can memorise noise or artifacts if optimisation proceeds too long, requiring ad hoc early stopping based on visual inspection wang2021early . Second, DIP’s pixel-wise reconstruction loss provides no explicit mechanism to control the frequency content or semantic structure of the recovered image. As a result, it tends to favor smooth, low-frequency components early in optimisation (due to spectral bias), while high-frequency details—such as texture or fine edges—are either fit too late or overfit entirely if optimisation proceeds unchecked rahaman2019spectral ; ulyanov2018deep .
@@
To mitigate these issues, recent work has explored two main directions cheng2019bayesian ; jo2021rethinking ; chen2020dip ; zhang2025survey . One line of research modifies the DIP architecture or optimisation dynamics—e.g., by freezing weights and optimising the input sun2021plug . Another line of work augments DIP with explicit regularisers, such as Total Variation (TV) rudin1992nonlinear ; liu2019image or Regularisation by Denoising (RED) romano2017little ; mataev2019deepred . From the perspective of inverse problems, such double-prior formulations are theoretically beneficial, as additional constraints help mitigate the inherent ill-posedness. However, these approaches fundamentally alter the nature of DIP by introducing explicit handcrafted or learned priors, additional hyperparameters, and tuning requirements. They move away from DIP’s core appeal: solving inverse problems using only the implicit regularisation of a randomly-initialised network, without any external data or assumptions. Once explicit regularisation is introduced, it becomes unclear whether performance gains stem from the network’s inductive bias or the added constraint. As such, direct comparisons with these methods obscure the role of implicit regularisation alone in reconstruction performance. In contrast, our goal is to retain the original spirit of DIP while addressing its limitations from within the single-prior framework.
@@
A new spectral-domain formulation of that disentangles frequency content. We reformulate image reconstruction as direct optimisation over Fourier coefficients, allowing fine-grained control over frequency modes. Unlike DIP, DSP enables structured frequency modulation and corrects DIP’s spectral bias. DSP imposes a spectral inductive bias aligned with both the statistics of images and the known learning dynamics of convolutional networks.
We establish the theoretical foundations of DSP and characterise its implicit regularisation:
– We prove that DSP biases reconstruction toward low-frequency structure while attenuating high-frequency noise, eliminating the need for early stopping. This is formalised through Theorem A.1, which characterises frequency-dependent convergence of the reconstruction error.
– We show that the DSP loss is continuously differentiable, locally Lipschitz, and supports stable descent under first-order optimisers (Theorem A.2). We further show that DSP exhibits spectrally ordered convergence aligned with the spectral bias of CNNs (Proposition A.1), explaining its robustness and stability across tasks.
–We show that DSP implicitly projects reconstructions onto a frequency-consistent manifold (Proposition A.3), achieves a favourable bias–variance tradeoff in the Fourier domain (Proposition A.3), and suppresses high-frequency noise without explicit priors (Proposition A.3), explaining its stability and performance across diverse inverse problems.
@@ 2 Methodology: Deep Spectral Prior
We introduce Deep Spectral Prior (DSP), a new formulation of Deep Image Prior (DIP) that shifts the reconstruction objective from pixel space to the frequency domain. While DIP minimises a pixel-wise loss, conflating frequency content and requiring early stopping, DSP directly matches Fourier coefficients, introducing a frequency-aware inductive bias. This leverages both the spectral structure of images and the spectral bias of CNNs, yielding sharper, more stable reconstructions without explicit regularisation. We show that this shift gives rise to distinct theoretical properties across four dimensions: regularisation, optimisation, geometry, and robustness.
@@ 2.1 From Pixels to Frequencies: Deep Spectral Prior
Let the observed degraded image be $y=Ax+\epsilon$, where $A$ is a known degradation operator and $\epsilon$ denotes additive noise. Let $\hat{x}_{\theta}=f_{\theta}(z)\in\mathbb{R}^{H\times W}$ denote the output of a convolutional network with fixed random input $z$, and let $\mathcal{F}(\cdot)$ denote the two-dimensional discrete Fourier transform (DFT).
Rather than minimising a spatial-domain loss, Deep Spectral Prior directly aligns Fourier coefficients via $\mathcal{L}(\theta):=\left\|\mathcal{F}(A\hat{x}_{\theta})-\mathcal{F}(y)\right\|_{2}^{2}.$ This reframes the objective from local pixel fidelity to global spectral coherence, reinforcing the natural low-frequency bias of images and CNNs. By decoupling signal from pixel-wise noise, DSP improves robustness and convergence. In the following, we show that DSP acts as an implicit spectral regulariser, aligned with both image statistics and network dynamics. We begin with a theoretical result that formalises how DSP naturally regularises reconstructions in the frequency domain. This reveals that DSP inherits the implicit spectral bias of CNNs but enhances it with a loss that prioritises low-frequency alignment.
@@
Deep Spectral Prior Loss. To incorporate inductive bias from the frequency domain, DSP modifies the reconstruction loss to operate in the Fourier domain. Specifically, the network parameters are optimised as:
$\theta^{*}=\arg\min_{\theta}\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2},$
where $\mathcal{F}$ denotes the (possibly non-normalised) discrete Fourier transform, $\mathcal{A}$ is the degradation operator, and $y$ is the observed corrupted image. This retains the self-supervised nature of DIP by relying only on the observed measurement $y$, while leveraging frequency information to guide reconstruction.
2.2 Spectral Regularisation and Implicit Bias of Deep Spectral Prior
We now formalise how the frequency-domain loss in DSP implicitly regularises the reconstruction. The following theorem shows that, under mild assumptions, DSP preserves low-frequency structure while attenuating high-frequency noise—providing a natural form of spectral denoising. This reveals that DSP inherits the implicit spectral bias of CNNs but enhances it with a loss that prioritises low-frequency alignment.
@@ 2.2 Spectral Regularisation and Implicit Bias of Deep Spectral Prior
(Spectral Regularisation in DSP).  Let $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be a convolutional neural network with parameters $\theta$, and let $z\in\mathbb{R}^{d}$ be a fixed random input. Let $y\in\mathbb{R}^{m}$ be a corrupted observation of an unknown clean image $x\in\mathbb{R}^{n}$, with a known corruption operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$. Let $\mathcal{F}:\mathbb{R}^{m}\rightarrow\mathbb{C}^{m}$ denote the discrete Fourier transform. Define the frequency loss: $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ Assume:
  1. 1.
    The network $f_{\theta}$ is trained using a smooth optimiser (e.g., gradient descent or Adam with small learning rate);
  2. 2.
    The network exhibits spectral bias, i.e., for training time $t$,
    $\left\|\frac{\partial}{\partial t}\mathcal{F}(f_{\theta}(z))(\omega)\right\|\text{ decreases as }\|\omega\|\text{ increases};$
  3. 3.
    The corruption noise $\eta=y-\mathcal{A}x$ is high-frequency, i.e.,
    $|\mathcal{F}(\eta)(\omega)|\text{ is large for high }\|\omega\|\text{ and small for low }\|\omega\|.$
Then the optimiser converges to $\hat{\theta}=\arg\min_{\theta}\mathcal{L}_{\mathrm{DSP}}(\theta)$ such that the reconstruction $\hat{x}=f_{\hat{\theta}}(z)$ satisfies:
$|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|\approx|\mathcal{F}(y)(\omega)|\text{for low }\|\omega\|,\quad\text{and}\quad|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|<|\mathcal{F}(y)(\omega)|\text{ for high }\|\omega\|.$
Hence, DSP suppresses high-frequency noise and does not require early stopping. (See proof in Appendix.)
To our knowledge, this is the first theoretical justification that using a Fourier magnitude loss (as in DSP) removes the need for early stopping in Deep Image Prior training. The result builds on spectral bias rahaman2019spectral properties but extends them with a formal optimisation perspective. Moreover, Theorem A.1 illustrates how DSP leverages both the spectral properties of images and the inherent frequency bias of neural networks. Low-frequency components, which dominate perceptual quality, are reconstructed accurately, while high-frequency noise—typically introduced by the degradation process—is suppressed. Crucially, this behaviour emerges without the need for early stopping, which is essential in classical DIP.
That is, Theorem A.1 establishes that the DSP loss, by prioritising low-frequency fidelity, inherently suppresses high-frequency noise—eliminating the need for early stopping, a crucial drawback in classical DIP. However, this result characterises the behaviour only at convergence. To gain a deeper understanding of how this behaviour emerges during training, we now turn to the optimisation trajectory of DSP. In particular, we ask: How does the reconstruction evolve in the frequency domain over time? The following proposition addresses this question by describing the frequency-wise convergence dynamics of the network output. It reveals that DSP not only converges to frequency-consistent reconstructions but does so in a spectrally ordered fashion, aligning with the spectral bias of convolutional networks.

### v2

@@ Abstract
We introduce the Deep Spectral Prior (DSP), a new framework for unsupervised image reconstruction that operates entirely in the complex frequency domain. Unlike the Deep Image Prior (DIP), which optimises pixel-level errors and is highly sensitive to overfitting, DSP performs joint learning of amplitude and phase to capture the full spectral structure of images. We derive a rigorous theoretical characterisation of DSP’s optimisation dynamics, proving that it follows frequency-dependent descent trajectories that separate informative low-frequency modes from stochastic high-frequency noise. This spectral mode separation explains DSP’s self-regularising behaviour and, for the first time, formally establishes the elimination of DIP’s major limitation—its reliance on manual early stopping. Moreover, DSP induces an implicit projection onto a frequency-consistent manifold, ensuring convergence to stable, physically plausible reconstructions without explicit priors or supervision. Extensive experiments on denoising, inpainting, and deblurring demonstrate that DSP consistently surpasses DIP and other unsupervised baselines, achieving superior fidelity, robustness, and theoretical interpretability within a unified, unsupervised data-free framework.
@@ 1 Introduction
Inverse problems lie at the heart of image restoration and reconstruction, encompassing tasks such as denoising, inpainting, super-resolution, and deblurring. Traditionally approached through handcrafted priors and variational optimisation [rudin1992nonlinear, chan2001active], these problems have seen a paradigm shift with the advent of deep learning, with single-image techniques effectively leverage data [essakine2025where, cheng2024singleshot, cheng2024you]. Among the most intriguing developments in this space is the Deep Image Prior (DIP) [ulyanov2018deep], which showed that the structure of an untrained convolutional neural network (CNN) alone can act as a powerful prior. By optimising a randomly initialised network to match a single degraded observation, DIP enables high-quality reconstructions—without requiring any training data. This surprising result has made DIP a widely adopted framework across diverse domains, including hyperspectral [rasti2021undip, li2025enhanced], microscopy [vu2021deep, zhou2020diffraction], compressed sensing [van2018compressed], and medical imaging [gong2018pet, yoo2021time, barbano2022educated, liang2025analysis, liang2025sequential].
@@
Despite its empirical success, DIP suffers from two key limitations that hinder its practical effectiveness. First, DIP is notoriously prone to overfitting. Due to the lack of explicit regularisation [rudin1992nonlinear, liu2019image, chang2000adaptive, chambolle2004algorithm, getreuer2012rudin], the network can memorise noise or artifacts if optimisation proceeds too long, requiring ad hoc early stopping based on visual inspection [wang2021early]. Second, DIP’s pixel-wise reconstruction loss provides no explicit mechanism to control the frequency content or semantic structure of the recovered image. As a result, it tends to favor smooth, low-frequency components early in optimisation (due to spectral bias [rahaman2019spectral]), while high-frequency details—such as texture or fine edges—are either fit too late or overfit entirely if optimisation proceeds unchecked [rahaman2019spectral, ulyanov2018deep].
@@
To mitigate these issues, recent works have explored two main directions [cheng2019bayesian, jo2021rethinking, chen2020dip, cascarano2022first, zhang2025survey]. One line of research modifies the DIP architecture or optimisation dynamics—e.g., by freezing weights and optimising the input [sun2021plug]. Another line of work augments DIP with explicit regularisers, such as Total Variation (TV) and variants [rudin1992nonlinear, liu2019image, chambolle2004algorithm, getreuer2012rudin] or Regularisation by Denoising (RED) [romano2017little, mataev2019deepred]. From the perspective of inverse problems, such double-prior formulations are theoretically beneficial [arndt2022regularization], as additional constraints help mitigate the inherent ill-posedness. However, these approaches fundamentally alter the nature of DIP by introducing explicit handcrafted or learned priors, additional hyperparameters, and tuning requirements. They move away from DIP’s core appeal: solving inverse problems using only the implicit regularisation of a randomly-initialised network, without any external data or assumptions. Once explicit regularisation is introduced, it becomes unclear whether performance gains stem from the network’s inductive bias or the added constraint. As such, direct comparisons with these methods obscure the role of implicit regularisation alone in reconstruction performance. In contrast, we propose the Deep Spectral Prior (DSP)—a novel spectral formulation that preserves its data-free nature while providing theoretical guarantees for stability and convergence.
@@
A new frequency-domain formulation. We introduce inverse imaging as spectral alignment in the Fourier domain, where optimisation operates directly on the amplitude and phase components of the signal. This formulation (Definition 1) enables explicit control over frequency modes and corrects the low-frequency bias inherent in DIP, yielding an intrinsic spectral inductive bias aligned with both image statistics and neural spectral dynamics.
We provide, for the first time, a rigorous analysis of the optimisation dynamics of the Deep Spectral Prior (DSP). This analysis characterises DSP’s implicit spectral regularisation through the following key results:
Spectral equivalence: We prove that the DSP loss is mathematically equivalent to the pixel-domain reconstruction error under a unitary Fourier transform (Corollary 2.1 ), preserving the reconstruction objective while reformulating the optimisation landscape.
Complex descent dynamics: We establish that DSP and DIP follow fundamentally different optimisation trajectories due to their underlying complex-valued descent laws (Theorem 2.1), providing a theoretical basis for DSP’s smoother and more stable convergence.
Spectral stability and early-stopping robustness: We show that our Deep Spectral Prior exhibits frequency-ordered convergence—recovering low-frequency components first while naturally suppressing high-frequency noise—thereby eliminating the need for early stopping (see Theorem 2.2 and Proposition 2.1).
@@ 2 Methodology: Deep Spectral Prior
Conventional pixel-space approaches such as the Deep Image Prior (DIP) optimise over spatial intensities, neglecting the spectral organisation that fundamentally governs natural images. We introduce the Deep Spectral Prior (DSP)—the first principled framework that reformulates image reconstruction entirely in the frequency domain and provides a theoretical understanding of its dynamics. By operating on complex-valued representations, DSP aligns learning with the intrinsic spectral bias of neural networks, offering provable stability, noise robustness, and self-regularising behaviour absent in spatial formulations.
@@ 2.1 From Pixels to Frequencies: Deep Spectral Prior
Let the observed degraded image be $y=\mathcal{A}x+\epsilon$, where $\mathcal{A}$ is a known linear degradation operator and $\epsilon$ denotes additive noise. Existing reconstruction approaches, including DIP, optimise pixel-wise fidelity but overlook global spectral coherence. Yet neural representations are inherently spectral, dominated by smooth low-frequency components and rapidly decaying high-frequency content.
This motivates a reformulation in the Fourier domain, where optimisation aligns frequency content rather than pixel intensities. Such a viewpoint naturally separates structured signal from stochastic noise and introduces frequency-dependent inductive bias with explicit control over spectral contributions. We next introduce the Deep Spectral Prior (DSP), a framework that performs reconstruction directly in the frequency domain, revealing a new form of implicit spectral regularisation (see Figure 1).
@@
Deep Spectral Prior (DSP). The Deep Spectral Prior (DSP) formulates image reconstruction directly in the frequency domain using complex-valued optimisation. Let $f_{\theta}(z)\in\mathbb{C}^{n\times n}$ denote the output of a neural network parameterised by weights $\theta$ for a fixed latent input $z\in\mathbb{C}^{d}$, and let $\mathcal{A}:\mathbb{R}^{n\times n}\!\to\!\mathbb{R}^{n\times n}$ be a known linear degradation operator acting on real-valued images. Given the observed corrupted image $y\in\mathbb{R}^{n\times n}$, the DSP loss is defined as
$\mathcal{L}_{\mathrm{DSP}}(\theta)$ $=\frac{1}{2}\big\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big\|_{2}^{2},$ (1)
$\theta^{*}$ $=\arg\min_{\theta}\mathcal{L}_{\mathrm{DSP}}(\theta),$
where $\mathcal{F}:\mathbb{R}^{n\times n}\!\to\!\mathbb{C}^{n\times n}$ is the discrete Fourier transform (DFT). All optimisation is performed in the complex domain, enabling the model to jointly learn amplitude and phase information and to capture the spectral correlations.
2.2 Spectral Equivalence and the Complex Descent Law of DSP
The Deep Spectral Prior (DSP) can be interpreted as a reformulation of image reconstruction in the frequency domain rather than the pixel domain. Although this shift might appear purely representational, it fundamentally alters the geometry of optimisation. We begin by establishing that the DSP loss is mathematically equivalent to the conventional pixel-space error, thereby preserving the reconstruction objective while transforming the optimisation landscape.
Figure 1: Optimisation trajectories for DIP, DSP, and conventional priors. DIP (red) overshoots and needs manual early stopping; the conventional prior (green) converges to a biased solution; DSP (blue) follows a stable spectral path and converges close to the GT without early stopping.
Let $x,\hat{x}\in\mathbb{R}^{m}$, and let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{C}^{m}$ denote the unitary discrete Fourier transform. Then
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}.$
In particular, for any neural network output $f_{\theta}(z)$ and observation $y$, $\mathcal{L}_{\mathrm{DSP}}(\theta)=\frac{1}{2}\big\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big\|_{2}^{2}$ is exactly equal to the pixel-domain reconstruction error $\frac{1}{2}\|\mathcal{A}f_{\theta}(z)-y\|_{2}^{2}$. Thus, DSP does not alter the objective value but reformulates the optimisation in the spectral domain. (See Proof in the Supplementary Material)
While the loss values in pixel and Fourier domains are equivalent under a unitary transform, the optimisation trajectories are not. In particular, real-valued networks, including that of DIP, and complex-valued networks (DSP) follow fundamentally different descent dynamics, as formalised below.
@@ 2.2 Spectral Equivalence and the Complex Descent Law of DSP
(Complex Descent Law: Distinct Optimisation Paths of DIP and DSP).  Consider the frequency-domain loss
$\mathcal{L}_{\mathrm{DSP}}(\theta)$ $=\frac{1}{2}\big\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big\|_{2}^{2}$ (2)
$=\frac{1}{2}\big(\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big)^{\dagger}\big(\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big).$
where $\mathcal{F}$ denotes the discrete Fourier transform and the Hermitian transpose. The optimisation dynamics under gradient descent differ fundamentally between real-valued DIP and complex-valued DSP:
  • DIP Optimisation: for $\theta\in\mathbb{R}^{p}$,
    $\frac{d\theta}{dt}=-\alpha\nabla_{\theta}\mathcal{L}_{\mathrm{DIP}}(\theta)$ (3)
    $=-\alpha\,\nabla_{\theta}\big(\mathcal{A}f_{\theta}(z)-y\big)^{T}\!\cdot\!\big(\mathcal{A}f_{\theta}(z)-y\big),$
    $\text{where}\,\,\mathcal{L}_{\mathrm{DIP}}(\theta)=\frac{1}{2}\big\|\mathcal{A}f_{\theta}(z)-y\big\|_{2}^{2}.$
  • DSP Optimisation: for $\tilde{\theta}=a+ib\in\mathbb{C}^{p}$,
[…]
    $=-\alpha\,\nabla_{\tilde{\theta}^{*}}\Big[\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)^{\dagger}\!\cdot\!\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)\Big].$
    where $\nabla_{\tilde{\theta}^{*}}$ denotes the Wirtinger gradient with respect to the complex conjugate variable. (See Proof in the Supplementary Material)
We undeline that our DSP formulation also applies when the network parameters are real-valued, since the real gradient corresponds to the real part of the Wirtinger derivative. Hence, DSP can be implemented with standard real-valued backbones, as the induced directional decomposition still captures the distinct curvature and convergence geometry of the spectral optimisation landscape.
2.3 The Spectral Stability Law: Robust Early-Stopping Dynamics of DSP
We next show that the complex-valued Deep Spectral Prior (DSP) naturally provides a more stable and extended early stopping window than the conventional Deep Image Prior (DIP). This property can be interpreted through the Neural Tangent Kernel (NTK) [jacot2018neural] framework, revealing that DSP induces smoother optimisation dynamics and mitigates premature overfitting.
Theorem 2.2.
(Robust Stopping Property of DSP).  For a DSP optimisation trajectory, the NTK approximation takes the form
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))=\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}$, (5)
[…]
where $\alpha$ is the gradient descent step size. The evolution of the discrepancy along gradient descent follows a linear ordinary differential equation (ODE) governed by the Hermitian operator $\tilde{\Phi}\tilde{\Phi}^{\dagger}$:
[…]
The ODE admits the closed-form solution
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}^{(t)}}(z))=\sum_{k=1}^{N}(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle\tilde{\psi}_{k},$ (8)
where $\tilde{\psi}_{k}$ and $\mu_{k}$ satisfy the eigendecomposition $\tilde{\Phi}\tilde{\Phi}^{\dagger}=\sum_{k=1}^{N}\mu_{k}\tilde{\psi}_{k}\tilde{\psi}_{k}^{\dagger}$. The mean-squared error (MSE) along the optimisation path of DSP evolves as
$MSE_{t}=\mathbb{E}\|\mathcal{F}(\mathcal{A}f_{\tilde{\theta}^{(t)}}(z))-\tilde{x}\|^{2}$ (9)
$=\sum_{k=1}^{N}\!\Big[(1-e^{-\alpha\mu_{k}t})^{2}(\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2})$
$-2(1-e^{-\alpha\mu_{k}t})Re\big\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\big\}+|\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}\Big]\!.$
where $\tilde{x}=\mathcal{F}(x)$, $\tilde{\mathcal{A}x}=\mathcal{F}(\mathcal{A}x)$ and noise $\tilde{\epsilon}=\mathcal{F}(\epsilon)$. The optimal stopping time for each mode $k$ is then
$t_{k}^{*}$ $=\frac{1}{\alpha\mu_{k}}\ln\!\left(\rho_{k}\right),$ (10)
$\rho_{k}$ $=\frac{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}}{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}-Re\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\}}.$
Different eigenvalues $\mu_{k}$ induce distinct convergence speeds, endowing DSP with an inherent robustness to stopping time. (See proof in the Supplementary Material)
Sketch of Proof. The NTK form in (28) arises from a first-order Taylor expansion of $\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))$ around $\tilde{\theta}_{0}$, assuming $\tilde{\Phi}(z)$ remains approximately constant in early training. This leads directly to the parameter-space ODE (29). Multiplying both sides by $\tilde{\Phi}(z)$ gives the discrepancy dynamics (30). Because $\tilde{\Phi}\tilde{\Phi}^{\dagger}$ is Hermitian and positive semidefinite, it admits the eigendecomposition $\tilde{\Phi}\tilde{\Phi}^{\dagger}=\sum_{k}\mu_{k}\tilde{\psi}_{k}\tilde{\psi}_{k}^{\dagger}$, which decouples the ODE into $N$ independent spectral components. Each mode evolves as $\frac{d\langle\tilde{\Phi}\tilde{u}-\tilde{y},\tilde{\psi}_{k}\rangle}{dt}=-\alpha\mu_{k}\langle\tilde{\Phi}\tilde{u}-\tilde{y},\tilde{\psi}_{k}\rangle$, yielding the closed-form solution (31).
Substituting into the MSE definition and using $\tilde{y}=\tilde{\mathcal{A}x}+\tilde{\epsilon}$ with $\mathbb{E}[\tilde{\epsilon}]=0$ removes cross-terms, resulting in (32). Differentiating $\mathrm{MSE}_{t}$ and solving $\frac{d\mathrm{MSE}_{t}}{dt}=0$ produces (33). Large $\mu_{k}$ correspond to smooth, low-frequency modes that converge rapidly, while small $\mu_{k}$ represent noise-dominated components that decay slowly. This spectral separation ensures stable convergence and naturally eliminates the need for explicit early stopping. (See full Proof in the Supplementary Material). $\square$
Figure 3: Blind denoising comparison on the ‘Plane’ image: our DSP method vs. unsupervised baselines, with zoomed-in views highlighting the retention of fine details.
The expression for $t_{k}^{*}$ reveals that modes associated with larger eigenvalues $\mu_{k}$ converge faster and capture the structured, low-frequency content of the image, while modes with small $\mu_{k}$ correspond to high-frequency noise. This spectral separation induces a stabilisation of optimisation: informative components settle early, whereas noise remains suppressed. Consequently, DSP self-regularises the training dynamics, removing the need for explicit early stopping. A higher-dimensional latent space further accentuates this frequency separation, enhancing convergence robustness.
This result provides, to our knowledge, the first theoretical justification showing that a Fourier-domain formulation—as introduced in DSP—inherently mitigates the need for early stopping. By optimising directly in the frequency domain, DSP prioritises low-frequency fidelity while attenuating slow-evolving high-frequency components, leading to stable convergence without manual intervention. Having established the asymptotic convergence behaviour, we next examine how this frequency selectivity emerges during training. In particular, we analyse the evolution of reconstruction in the spectral domain to characterise the implicit frequency ordering of DSP optimisation.